# An Efficient Learning Control Framework with Sim-to-Real for String-Type Artificial Muscle-Driven Robotic Systems

### v3

@@ V-C Bootstrap via PID
On-line DRL agents primarily rely on random exploration [24]. However, due to the slow response characteristics of artificial muscles, random exploration policies often fail to explore the state space comprehensively and efficiently, leading to low sample efficiency. To address this limitation, we accelerate the early-stage training by incorporating demonstration data into the initial replay buffer. For muscle-driven robotic systems, PID controllers are typically sufficient to approach target states and deliver moderate control performance. As such, they can generate trajectories that explore the defined state space more comprehensively, providing a beneficial starting point for the DRL learning process.
Specifically, we initialize a PID controller that offers a barely acceptable (and typically not excellent) control performance before DRL training begins. Within the first $M$ episodes of training, the PID controller interacts with the environment and generates trajectories, which are then stored in the replay buffer along with the rewards. Unlike traditional behavior cloning methods, we do not employ a supervised loss for demonstration learning, allowing the agent to avoid strictly mimicking the PID control policy. Subsequently, the agent continues interacting with the environment to refine its policy, eventually surpassing the PID controller and achieving superior control performance.
@@ V-D Augmentation of State Vector
Data augmentation refers to a class of techniques aimed at increasing the size and improving the quality of training datasets, thereby facilitating the development of more robust deep learning models[42]. The fundamental concept of data augmentation involves transforming limited data into diverse and functionally equivalent datasets. Given the labor-intensive and computationally expensive nature of collecting training data in robotic systems, we integrate data augmentation into the agent’s learning process. This paper proposes a novel and straightforward state-based data augmentation technique designed for our targeted robotic systems.
@@
As commonly acknowledged, rewards serve as pivotal guides in the learning process of DRL agents. Therefore, our objective is to generate additional data transitions with different rewards. As described in Section III, the reward $r_{t}$ is calculated as:
@@
which creates different transition tuples at a negligible computational cost.
In this paper, we denote the bootstrap method by the letter ’B’, augmentation method by the letter ’A’, and the randomization of muscle dynamics by the letter ’R’. Thus, our proposed SAC algorithm with all three improvements is denoted by SAC-BAR. The pseudo code of our overall SAC-BAR learning control framework is presented in Algorithm 1.
Figure 6: Reward trajectories with respect to the training episode number. (a) Comparison results between our SAC-BAR and SOTA methods including REDQ, DroQ, RLPD, and X-QL. (b) Results of ablation study where ‘B’, ‘A’, and ‘R’ indicate bootstrap, augmentation, and randomization techniques, respectively.
@@ VI Simulation Experiments
In this section, we present the training results from simulation environments, comparing our proposed SAC-BAR against several SOTA reinforcement learning algorithms in terms of sample efficiency on two robotic systems. Additionally, we perform ablation studies to evaluate the effectiveness of the proposed bootstrap and augmentation methods.
@@ VI-A Training Setup
At the beginning of each training episode, a target state is randomly selected from a uniformly distributed range. For instance, in the robotic eye scenario, the target angles adhere to $\theta^{*}\sim\mathcal{U}[-15^{\circ},15^{\circ}]$ and $\psi^{*}\sim\mathcal{U}[-10^{\circ},10^{\circ}]$. The robotic agent strives to reach the target from the initial state throughout the episode. In this paper, simulation environments are all written in Python. Training and testing in simulation are conducted on a computer with an Intel Xeon Platinum 8383C CPU and an NVIDIA GeForce RTX 4090 GPU. The algorithms are implemented in PyTorch and trained using the Adam optimizer. The learning rate of DRL algorithms is set to $3\times 10^{-4}$ and the hidden layer size is fixed at $256$. The training hyperparameters $N$ and $M$ are empirically chosen based on the system’s complexity. Specifically, $N=2000$ and $M=250$ are used for the robotic eye, while $N=3500$ and $M=500$ are used for the robotic wrist. When the bootstrap method is not applied, random exploration is employed during the initial $N$ warm-up episodes. These hyperparameters are kept consistent across all comparison and ablation studies.
@@ VI-B Comparison Results Against SOTA Methods
Several SOTA sample-efficient RL algorithms, designed for state-based observations and continuous action spaces, are selected for comparison. REDQ[34] integrates an ensemble of Q-functions with in-target minimization across a random subset of these functions. This allows for training with a high Update-To-Data (UTD) ratio, significantly improving sample efficiency. DroQ[35] equips the Q-functions with dropout connections and layer normalization techniques, maintaining a high UTD ratio. RLPD[36] leverages offline data during online learning by loading demonstrations into the replay buffer and oversampling them during training. X-QL[37] directly models soft optimal values and introduces a novel optimization framework based on the Gumbel distribution.
We trained the proposed SAC-BAR on two robotic systems and compared its performance with these SOTA approaches. Performance was measured by the online per-step average reward achieved by the agents, with each algorithm trained across four independent runs. Fig. 6(a) presents the comparative results between SAC-BAR and the selected SOTA RL methods. In the robotic eye system, SAC-BAR achieves performance comparable to (and slightly better than) DroQ and RLPD, while significantly outperforming REDQ and X-QL in terms of sample efficiency. In the robotic wrist system, SAC-BAR not only significantly outperforms RLPD and DroQ but also demonstrates moderate improvements in convergence speed over REDQ and X-QL. Notably, REDQ’s performance relies on a high UTD ratio and an ensemble of Q-functions, resulting in high computational cost; for instance, REDQ requires 18.2 hours for 3,500 training episodes, whereas SAC-BAR completes the same training in only 4.1 hours under our experimental setup. Moreover, X-QL exhibited significant instability during the early stages of training in both robotic systems. Overall, these results demonstrate that SAC-BAR is better suited than the existing SOTA RL methods for the string-type artificial muscle robots targeted in this study.
@@ VI-C Ablation Study in Simulation
To evaluate the effectiveness of the proposed bootstrap and augmentation techniques, an ablation study was conducted on two robotic systems. Three variants of our proposed framework were tested as follows: (1) w/o A: This variant excludes the state vector augmentation method. (2) w/o B: This variant removes the bootstrap via PID method. (3) w/o AB: This baseline variant excludes both mechanisms. Each algorithm was trained and evaluated in simulation under randomized dynamics across four independent runs.
Training results of our proposed SAC-BAR, with and without bootstrap and augmentation, are presented in Fig. 6(b). As illustrated in the figures, our proposed algorithm exhibits the highest level of data efficiency in both robotic systems. Compared to the SAC algorithm trained solely with dynamics randomization (referred to as “w/o AB”), the integration of bootstrap and augmentation enhances data efficiency and expedites the training process. In the robotic eye system, the baseline algorithm achieves an average reward of $-0.10$ after 2,000 training episodes, whereas our approach achieves comparable performance in only approximately 750 episodes, resulting in a 2.67-fold increase in data efficiency compared to the original level. Similarly, in the robotic wrist system, the baseline algorithm achieves an average reward of $-1.61$ after 3,500 episodes, while ours achieves comparable performance in only about 1,200 episodes, leading to a 2.92-fold increase in data efficiency. Notably, in the case of the more complex robotic wrist system, we find the demonstrations provided by the PID controller yield a more significant acceleration effect during the early stages of training. Meanwhile, the augmentation method demonstrates a favorable impact on enhancing the training speed in both systems and throughout the entire learning process of the algorithms.
@@ VII Real-World Experiments
This section presents the real-world experimental results of the proposed SAC-BAR on two testing robotic platforms. We first demonstrate the results of an ablation study conducted on the robotic wrist platform, which validate the effectiveness of the proposed dynamics randomization in enhancing control performance in real-world scenarios. Next, experimental results on the robotic eye are provided to show the generalizability of SAC-BAR across similar string-type artificial muscle-driven robotic systems. Additionally, we investigate the impact of the randomization range on control performance.
VII-A Ablation Study on Robotic Wrist
Figure 7: Comparison of experimental trajectories between the baseline and SAC-BAR. Target angles are set at ($\theta^{*}=2.5^{\circ},\phi^{*}=-7.5^{\circ}$). (a), (b) The system output angles ($\theta,\phi$) and corresponding actions $a_{t}$ of the baseline SAC algorithm. (c), (d) Those of our proposed SAC-BAR algorithm.
To evaluate the impact of each component on real-world control performances, three different control policies are tested on the robotic wrist platform, including the vanilla SAC algorithm, SAC with only randomization of muscles (SAC-R), and our proposed SAC-BAR. The parallel robotic wrist starts from the initial state with $\theta=\phi=0^{\circ}$ and attempts to reach and maintain the target angle set $(\theta^{*},\phi^{*})$. The action vector $a$ is computed by the trained actors at a time step of $t_{a}=0.5$ s, and each episode spans a duration of $t_{\text{total}}=20$ s. In Fig. 7(a), it is evident that without randomization of muscle dynamics, the vanilla-SAC struggles to achieve the desired orientation. While the agent may eventually reach the desired orientation at certain target angles, maintaining that target angle becomes a challenge, resulting in a considerable steady-state error. Conversely, as observed in Fig. 7(c), the proposed SAC-BAR agent achieves the target state in less than $10$ s and consistently remains within its neighborhood thereafter. As illustrated by the input action trajectories in Fig. 7(d), this success is attributed to the agent learning to maximize the action value during the rising phase and promptly reduce it upon approaching the target angles. Conversely, in Fig. 7(b), there exist serious oscillations in the action trajectories of vanilla-SAC during the initial $10$ s, leading to a prolonged rise time and sometimes even failure to reach the target.
@@ VII-A Ablation Study on Robotic Wrist
In order to comprehensively evaluate the control performance of the algorithms across different target angles, we designed a steady-state error field test for the robotic wrist. We selected a series of target angles $(\theta^{*},\phi^{*})$ spaced at $2.5^{\circ}$ intervals within the defined range $\theta^{*},\phi^{*}\in[-10^{\circ},10^{\circ}]$. Each agent was set to run for $25$ s, sequentially targeting each angle pair. The steady-state error is measured and averaged over the last $5$s window, determined by computing the Euclidean distance norm between the current and target angles, defined as $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\phi-\phi^{*})^{2}}$. The steady-state tracking errors for different targets of vanilla-SAC, SAC-R, and SAC-BAR are presented in Fig. 8(a), (b), and (c), respectively. The $\mathrm{x}$ and $\mathrm{y}$ axes therein represent $\theta^{*}$ and $\phi^{*}$, while the color bars represent the averaged steady-state errors. Following that, box charts are illustrated in Fig. 8(d). These figures demonstrate that vanilla-SAC agent underperforms across nearly the entire defined range, whereas both SAC-R and SAC-BAR show significantly improved control performance.
The steady-state errors averaged over the defined range, denoted as $\bar{e}_{\text{ss}}$, for three variants and the PID controller are presented in Table III. In real-world experiments, all three DRL-based controllers demonstrate superior performance compared to the traditional PID controller, reducing the tracking error by $11.6\%$, $35.9\%$, and $43.7\%$, respectively. Compared to the baseline SAC, SAC-R and SAC-BAR achieved reductions in the averaged tracking errors of $27.9\%$ and $36.3\%$, respectively. Specifically, our proposed SAC-BAR reaches the desired orientation at every target angle pair with an averaged error of $0.58^{\circ}$. Furthermore, as shown in Table III, the observed improvement in real-world performance can be attributed to overcoming the sim-to-real gap, as evidenced by the reduced disparity between simulated and actual results.
TABLE III: The experimental (Real) and simulation (Sim) results of steady state error using different control methods
@@
PID SAC SAC-R SAC-BAR
Real $1.03\pm 0.49$ $0.91\pm 0.46$ $0.66\pm 0.39$ $0.58\pm 0.37$
Sim $1.12\pm 0.47$ $0.46\pm 0.27$ $0.47\pm 0.33$ $0.40\pm 0.17$
The above experimental results indicate that the proposed randomization of muscle dynamics significantly enhances the real-world control performance in the robotic systems of interest, while bootstrap and augmentation techniques contribute relatively less to the improvement. We believe that the enhancement observed in SAC-BAR compared to SAC-R primarily stems from the improvement in training efficiency. With the incorporation of bootstrap and augmentation methods, SAC-BAR achieves more comprehensive training and higher averaged rewards in simulation after 3,500 training episodes. We believe this improvement is potentially more significant in scenarios where exploration is more challenging.

### v4

@@ V-C Bootstrap via PID
On-line DRL agents primarily rely on random exploration [21]. However, due to the slow response characteristics of artificial muscles, random exploration policies often fail to explore the state space comprehensively and efficiently, leading to low sample efficiency. To address this limitation, we accelerate the early-stage training by incorporating demonstration data into the initial replay buffer. For muscle-driven robotic systems, PID controllers are typically sufficient to approach target states and deliver moderate control performance. As such, they can generate trajectories that explore the defined state space more comprehensively, providing a beneficial starting point for the DRL learning process. Specifically, we initialize a PID controller that offers a barely acceptable (and typically not excellent) control performance before DRL training begins. Within the first $M$ episodes of training, the PID controller interacts with the environment and generates trajectories, which are then stored in the replay buffer along with the rewards. Compared to using explicit pre-training with PID data or a supervised loss for demonstration learning, our design mitigates the potential risk of the policy overfitting to suboptimal PID behavior before transitioning to RL.
@@ V-D Augmentation of State Vector
Data augmentation refers to a class of techniques aimed at increasing the size and improving the quality of training datasets[41]. The concept involves transforming limited data into diverse and functionally equivalent datasets. Given the labor-intensive and computationally expensive nature of collecting training data in robotic systems, we integrate data augmentation into the agent’s learning process. This paper proposes a novel and straightforward state-based data augmentation technique designed for our targeted robotic systems.
@@
Our objective is to generate additional data transitions with different rewards. As described in Section III, the reward $r_{t}$ is calculated as:
@@
which creates different transition tuples at a negligible computational cost. We choose to perturb the target vectors instead of the state vectors because perturbing state vectors may introduce physical inconsistencies, which can prevent the agent from learning accurate dynamics. The pseudo code of SAC-BAR learning control framework is presented in Algorithm 1.
Figure 6: Reward trajectories with respect to the training episode number. The shaded regions indicate the standard deviation of the rewards across multiple runs. (a) Comparative results between our proposed SAC-BAR method and SOTA approaches, including REDQ, DroQ, RLPD, and X-QL. (b) Comparative results from the ablation study, where ‘A’, ‘Ps’, ‘B’, and ‘Pt’ represent the augmentation method, state perturbation variation, bootstrap mechanism, and pre-training variation, respectively.
@@ VI Simulation Experiments
This section presents the training results from simulation environments, comparing our proposed SAC-BAR against several SOTA reinforcement learning algorithms in terms of sample efficiency on two robotic systems. Additionally, we perform ablation studies to evaluate the effectiveness of the proposed methods and discuss the choice of hyperparameters.
@@ VI-A Training Setup
In each training episode, the robotic agent aims to reach a target state randomly selected from a uniformly distributed range. The initial PID controller parameters are set as $K_{p}=2.1$, $K_{i}=0.2$, and $K_{d}=0.5$ for the robotic eye, and $K_{p}=3.3$, $K_{i}=0.5$, and $K_{d}=0.3$ for the robotic wrist. The number of data augmentation is set as $n=10$. The neural network architecture for both the actor and critic consists of a fully connected input layer, a GRU layer with 256 hidden units, and a fully connected output layer, with ReLU as the activation function. The model is trained using a batch size of 20 trajectories and an experience replay buffer with a capacity of $1\times 10^{5}$ trajectories. The learning rate is set to $3\times 10^{-4}$. Additionally, the SAC temperature parameter $\alpha$ is configured to be auto-tuned, as described in[26]. The hyperparameters $N$ and $M$ are empirically determined based on the system’s complexity. Specifically, $N=2000$ and $M=250$ are used for the robotic eye, while $N=3500$ and $M=500$ are used for the robotic wrist. When the bootstrap method is not applied, random exploration is employed during the initial $N$ warm-up episodes. The algorithms are implemented in PyTorch and trained on a computer with an Intel Xeon Platinum 8383C CPU and an NVIDIA GeForce RTX 4090 GPU.
@@ VI-B Comparison Results Against SOTA Methods
Several SOTA sample-efficient RL algorithms, tailored for state-based observations and continuous action spaces, are selected for comparison. These include REDQ[31], DroQ[32], RLPD[33], and X-QL[34], which represent diverse approaches to improving sample efficiency. We train SAC-BAR on two robotic systems and compare its performance with these SOTA approaches. Performance is measured by the online per-step average reward achieved by the agents, with each algorithm trained across four independent runs. Fig. 6(a) presents the comparative results between SAC-BAR and the selected SOTA RL methods. In the robotic eye system, SAC-BAR achieves performance comparable to (and slightly better than) DroQ and RLPD, while significantly outperforming REDQ and X-QL in terms of sample efficiency. In the robotic wrist system, SAC-BAR not only significantly outperforms RLPD and DroQ but also demonstrates moderate improvements in convergence speed over REDQ and X-QL. Notably, REDQ’s performance relies on a high UTD ratio and an ensemble of Q-functions, resulting in high computational cost; for instance, REDQ requires 18.2 hours for 3,500 training episodes, whereas SAC-BAR completes the same training in only 4.1 hours under our experimental setup. Moreover, X-QL exhibits significant instability during the early stages of training in both robotic systems. These results demonstrate that SAC-BAR is better suited than the existing SOTA methods for the string-type artificial muscle robots targeted in this study.
@@ VI-C Ablation Study in Simulation
To evaluate the effectiveness of the proposed bootstrap and augmentation techniques, an ablation study is conducted on two robotic systems. Five variants of our proposed framework are tested as follows: (1) Baseline: This variant employs only the SAC algorithm without any additional components. (2) w/ A: This variant incorporates the augmentation method. (3) w/ Ps: This variant perturbs the state vectors instead of target vectors. (4) w/ B: This variant integrates the bootstrap mechanism. (5) w/ Pt: This variant pre-trains the networks on data generated by the PID controller before exploration.
Training results of these variants and our proposed SAC-BAR are presented in Fig. 6(b). As illustrated in the figures, our proposed algorithm exhibits the highest level of data efficiency in both robotic systems. Compared to the baseline SAC algorithm, the integration of bootstrap and augmentation enhances data efficiency and expedites the training process. In the robotic eye system, the baseline algorithm achieves an average reward of $-0.10$ after $2,000$ training episodes, whereas our approach achieves comparable performance in only approximately $750$ episodes, resulting in a $2.67$-fold increase in data efficiency compared to the original level. Similarly, in the robotic wrist system, the baseline algorithm achieves an average reward of $-1.61$ after $3,500$ episodes, while ours achieves comparable performance in only about $1,200$ episodes, leading to a $2.92$-fold increase in data efficiency. The performance of the variant perturbing state vectors is lower compared to the baseline. We attribute this to the perturbation of state variables, which likely hinders the agent’s ability to accurately learn the system dynamics. While pre-training on PID-generated data enhances training efficiency, we observe that overfitting may occur during the pre-training phase, leading to a significant drop in the average return.
VI-D Influence of Hyperparameters
Figure 7: (a) Comparative results for different scaling factors of the initial PID gains ($\beta=0.5,1.0,2.0,3.0$), where $K^{\prime}_{p}=\beta K_{p}$, $K^{\prime}_{i}=\beta K_{i}$, and $K^{\prime}_{d}=\beta K_{d}$. (b) Comparative results for different numbers of augmentation samples ($n=1,5,10,20$).
The hyperparameters that significantly affect the performance of SAC-BAR include the initial PID gains and the number of augmentation samples. To evaluate their impact, we train the algorithm under different hyperparameter settings on the robotic wrist. Specifically, we compare the training results using different scaling factors for the initial PID gains ($\beta=0.5,1.0,2.0,3.0$) and varying numbers of augmentation samples ($n=1,5,10,20$), as shown in Fig. 7. The results indicate that our method is generally robust to these hyperparameters. When the PID gains are within a reasonable range (not necessarily optimal), the differences in training results are minor. However, when the PID gains are extreme (e.g., $\beta=3$), the demonstrations it provides are of poor quality, thus affecting the algorithm’s training performance. Similarly, the performance of the model trained with $n=20$ drops significantly, as excessive augmentation introduces redundancy, causing the agent to overfit to the augmented data rather than focusing on the true dynamics of the environment. On the other hand, an extremely low $n$ reduces the diversity of the training data, slightly degrading performance. A moderate value of $n$ (e.g., $n=5$ or $n=10$) strikes a balance between data diversity and training stability.
@@ VII Real-World Experiments
This section presents the real-world experimental results of the proposed SAC-BAR. We first demonstrate the results of an ablation study conducted on the robotic wrist platform, which validate the effectiveness of the proposed dynamics randomization in enhancing control performance in real-world scenarios. Next, experimental results on the robotic eye are provided to show the generalizability of SAC-BAR across similar robotic systems. Additionally, we investigate the impact of the randomization variance on control performance.
VII-A Comparison Results on Robotic Wrist
Figure 8: Comparison of experimental trajectories between the baseline and SAC-BAR. Target angles are set at ($\theta^{*}=2.5^{\circ},\phi^{*}=-7.5^{\circ}$). (a), (b) The system output angles ($\theta,\phi,e_{\text{ss}}$) and corresponding actions $a_{t}$ of the baseline SAC algorithm. (c), (d) Those of our proposed SAC-BAR algorithm.
To evaluate the impact of each component on real-world control performances, three different control policies are tested on the robotic wrist platform, including the vanilla SAC algorithm, SAC with only randomization of muscles (SAC-R), and our proposed SAC-BAR. The parallel robotic wrist starts from the initial state with $\theta=\phi=0^{\circ}$ and attempts to reach and maintain the target angle set $(\theta^{*},\phi^{*})$. The action vector $a$ is computed by the trained actors at a time step of $t_{a}=0.5$ s, and each episode spans a duration of $t_{\text{total}}=20$ s. In Fig. 8(a), it is evident that without randomization of muscle dynamics, the SAC agent struggles to achieve the desired orientation. While the agent may eventually reach the desired orientation at certain target angles, maintaining that target angle becomes a challenge, resulting in a considerable steady-state error. Conversely, as observed in Fig. 8(c), the proposed SAC-BAR agent achieves the target state in less than $10$ s and consistently remains within its neighborhood thereafter. As illustrated by the input action trajectories in Fig. 8(d), this success is attributed to the agent learning to maximize the action value during the rising phase and promptly reduce it upon approaching the target angles. Conversely, in Fig. 8(b), there exist serious oscillations in the action trajectories of SAC during the initial $10$ s, leading to a prolonged rise time and sometimes even failure to reach the target.
@@ VII-A Comparison Results on Robotic Wrist
To comprehensively evaluate the control performance of the algorithms across different target angles, we design a steady-state error field test for the robotic wrist. We select a series of target angles $(\theta^{*},\phi^{*})$ spaced at $2.5^{\circ}$ intervals within the defined range $\theta^{*},\phi^{*}\in[-10^{\circ},10^{\circ}]$. Each agent is set to run for $25$ s, sequentially targeting each angle pair. The steady-state error is measured and averaged over the last $5$s window, determined by computing the Euclidean distance norm between the current and target angles, defined as $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\phi-\phi^{*})^{2}}$. The steady-state errors for different targets of SAC, SAC-R, and SAC-BAR are presented in Fig. 9(a), (b), and (c), respectively. The $\mathrm{x}$ and $\mathrm{y}$ axes therein represent $\theta^{*}$ and $\phi^{*}$, while the color bars represent the averaged steady-state errors. Following that, box charts are illustrated in Fig. 9(d). These figures demonstrate that SAC agent underperforms across nearly the entire defined range, whereas both SAC-R and SAC-BAR show significantly improved control performance. These results indicate that the proposed randomization of muscle dynamics significantly enhances the real-world control performance.
TABLE III: The simulation (Sim) and experimental (Real) results of steady state error using different control methods
@@
PID AC PPO SAC SAC-R SAC-BAR
Sim 1.12±$\pm 0.47$ 0.80±$\pm 0.41$ 0.79±$\pm 0.32$ 0.46±$\pm 0.27$ 0.47±$\pm 0.33$ 0.40±$\pm 0.17$
Real 1.03±$\pm 0.49$ 0.86±$\pm 0.67$ 1.24±$\pm 0.48$ 0.91±$\pm 0.46$ 0.66±$\pm 0.39$ 0.58±$\pm 0.37$
To evaluate the real-world performance of our method against traditional control approaches, we implement the initial PID controller, an adaptive controller (AC), and the Proximal Policy Optimization (PPO) algorithm[42] for comparison. The steady-state errors averaged over the defined range, denoted as $\bar{e}_{\text{ss}}$, for these methods are tested in both simulation and real-world scenarios. The experimental results are summarized in Table III. In real-world experiments, our method achieves superior performance with a mean error of $0.58^{\circ}$, significantly lower than those of the other methods. Specifically, compared to PID, AC, PPO, and SAC, our method reduces the mean error by $43.7\%$, $32.6\%$, $53.2\%$, and $36.3\%$, respectively. While PPO and SAC are strong DRL baselines, they struggle with the sim-to-real gap, resulting in larger errors and less stable behaviors. The adaptive controller performs better due to its robustness to parameter uncertainties, but it requires extensive parameter tuning and still exhibits suboptimal performance at certain target angles. A comparison of SAC and SAC-R results demonstrates that the randomization of muscle dynamics effectively bridges the sim-to-real gap, leading to reduced disparity between simulated and actual results. Furthermore, SAC-BAR achieves more comprehensive training in simulation than SAC-R, resulting in a slight improvement in control performance.